# https://cs.uwaterloo.ca/~smwatt

### v1

@@ 7 The Role of Standards
  • the XML, MathML and OpenMath standards [51, 52, 6].
@@
In this spirit, the CSL implementation of Lisp, has provided all of the short floating point required by Common Lisp. The revision of Aldor, currently underway, is based on the C++20 standard, even though support for it is still only partial in the most popular C++ implementations.
@@
POSIX + Unicode + IEEE 754 floating point + gcc $\cap$ clang
@@
  • the sizes for integer arithmetic (16, 18, 32, 36, 64 bits) and consequently the transition to big integers,
@@
  • the size and representation of pointers (e.g. 16, 18, 20, 32, 36, … bits, segment-offset pair versus flat address space, etc)a,
@@
One consequence of the reasonable assumptions that can be made for modern platforms is that the run-time support for systems has some portable-in-practice options for inexpensive tagging data by pointer encoding. The two most common techniques involve placing information in the low order or high order bits of pointers. If a system allocates dynamic memory in multiples of $2^{n}$ bytes, then pointers to dynamically allocated objects will always have zero as their lower $n$ bits. This allows those bits to be used for other purposes, for example as type tags. Most widely used systems today use only the lower 48 bits of 64 bit pointers for virtual addresses, replicating the 48th bit for the higher bits. The adoption of IEEE 754 floating point numbers allows the 48 bits of pointers to be encoded as NANs, with additional bits available for tags. [^2] However, as larger virtual address spaces are supported, such as by Intel’s 5-level page tables, this becomes more complicated.
We anticipate that being standard compliant versus having a finite number of low-level specialisations will remain a nuanced judgement call. The issue of claimed versus actual compliance of language implementations will remain ongoing as standards are periodically revised. For example, the format library was initially unavailable in the most popular C++ imnplementations compiling to the C++20 standard. Said another way, claimed standard compliance remains untrustworthy.
@@
  • libraries of formalized mathematical proofs (e.g. Lean’s mathlib, Mizar, Isabelle/HOL),
@@
As yet, these areas are not problematic, but it is not unreasonable to anticipate that one or more of them may become difficulties until standardised.
@@ 8 Portability of Mathematics
The portability of mathematics itself remains an on-going open issue in the portability of computer algebra. Differing definitions occur across mathematics’ many fields. We give a few examples:
  • Logic and foundations Do the natural numbers include zero or not? Does the logic include the law of the excluded middle?
  • Algebra Rings may be or may not be required to have a multiplicative identity. Lie brackets may be defined as $[x,y]=u(xy-yx)$ with $u$ being $1$ or $i$. Modules are sometimes implicitly taken to be left modules, but in noncommutative contexts the distinction between left and right multiplication matters.
  • Algebraic geometry Do divisors on a scheme follow the Weil or Cartier definition? These notions coincide on smooth varieties, but do not on singular varieties.
  • Algebraic topology Poincaré duality is standard in oriented manifolds, but treatments differ in non-orientable cases.
  • Homological algebra Boundary operators in simplicial homology can differ by sign conventions, also affecting chain complexes.
  • Graph theory Do graphs allow loops, that is edges from a vertex to itself? Are complete graphs and edgeless graphs considered to be strongly regular? This affects their classification and algebraic combinatorics.
  • Complex functions Branch cuts can differ for multivalued functions. For example the principal value of complex logarithm may be defined as $\log z=\log|z|+i\arg z$, where $\arg z\in(-\pi,\pi]$ or $[0,2\pi)$, depending on convention. Definitions of Fourier transforms vary on whether the forward or inverse transform has the negated exponent and how the forward and inverse transforms are normalized.
  • Differential geometry The Riemann curvature tensor differs by a sign depending on convention. This affects derived quantities such as the Ricci tensor and scalar curvature. The Hodge star operator and volume forms depend on the metric signature and orientation.
@@
A second form of impediment to the portability of mathematics is the encoding of mathematical algorithms. Signifigant tracts of mathematical knowledge are captured in the syntax of computer algebra and proof assistant systems. Examples were given in Section 2 of two approaches to symbolic integration locked in unfrequented systems.
@@ 9 Final Thoughts and Conclusions
In the platform consolidation period, the set of suitable widely used environments simplified. Computer architectures were for the most part using 8-bit bytes, memory was byte addressable, and the Unicode, IEEE 754 and XML standards could be relied upon for the most part for data representation. The posix standard provided common abstractions across operating systems, though bespoke memory management was still widely used. Most systems or stand-alone programs were implemented in Lisp, C, C++ or Fortran, all of which had mature language standards with mostly compliant implemenations. Freely available cross-platform implementations meant implementors could rely on one particlar implementation (e.g. g++ or clang/llvm). It remains the case that older legacy systems or modern embedded systems may have diverse architectural considerations that can be handled by earlier methods.
We are now in the mathematical consolidation period where the main remaining portability issues are at the levels of mathematical definitions and algorithm encoding. These come into play when it is desired to share mathematical data or computations across different computer algebra systems and mathematical proof assistants. The issue of consistency of mathematical definitions has existed since well before the digital computer era. Seemingly slight differences in the meaning of specific mathematical symbols can make what is an identity or theorem in one system be false in another system. Examples of issues that arise include sign conventions, normalization factors, branch cut choices for inverse complex functions, and foundational logics. The issue of algorithm encoding is that much of mathematical knowledge is captured in the form of programs written in the application-level programming language of a handful of systems such as Magma, Maple, Mathematica, Reduce and SageMath (for computer algebra) or Mizar, Isabelle/HOL or Lean (for proof assistants). Here, not only may the details of definitions differ, but the encoding of algorithms is by-and-large completely incompatible.
@@
The second conclusion is that portability is not automatically easy, and the main challenges have altered over time-scales that are relevant to some software. As outlined above, we began in the 1960s with a small number of wildly different machines, but software that was simple by today’s standards. Over the next decades, into the 1990s, many new platforms were introduced and standardisation of languages and interfaces was a work in progress. We have reached apoint where there are fewer machines that are relevant again, and much standardisation to help, but software is much more ambitious. Additionally, systems are now concerned with security or protecting intellectual property so things are sometimes tough again!
@@
The fourth conclusion is that freely available algorithmic code that runs atop a commercial platform may feel locked down if it is complex. Arranging that it is portable to different worlds (either commercial or open source) is a long way from automatic. So the fact that somebody has a paper describing how they can do something you like may or may not help. And if they then cease to be available to support what they done the progress they made may be lost. So portability there is critical!
The final conclusion is that the people who care about portability are mostly invisible, but rest of the world depends on the work they do. Take, for example, the many many who keep Linux afloat. For (much) smaller projects there is a (much) smaller pool of expertise, but maintenance and porting are very close neighbours.

### v4

@@ 7 The Role of Standards
  • the XML, MathML and OpenMath standards [55, 56, 49, 9].
@@
In this spirit, the CSL implementation of Lisp, has provided all of the short floating point arithmetic required by Common Lisp. The revision of Aldor, currently underway, is based on the C++20 standard, even though support for it is still only partial in the most popular C++ implementations.
@@
Posix + Unicode + Ieee 754 floating point + gcc/clang
@@
  • the sizes for native integer arithmetic (16, 18, 32, 36, 64 bits) and consequently the transition to big integers,
@@
  • the size and representation of pointers (e.g. 16, 18, 20, 32, 36, … bits, segment-offset pair versus flat address space, etc),
@@
One consequence of the reasonable assumptions that can be made for modern platforms is that the run-time support for systems has some portable-in-practice options for inexpensive tagging data by pointer encoding. The two most common techniques involve placing information in the low-order or high-order bits of pointers. If a system allocates dynamic memory in multiples of $2^{n}$ bytes, then pointers to dynamically allocated objects will always have zero as their lower $n$ bits. This allows those bits to be used for other purposes, for example as type tags. Most widely used systems today use only the lower 48 bits of 64 bit pointers for virtual addresses, replicating the 48th bit for the higher bits. The adoption of Ieee 754 floating point numbers allows the 48 bits of pointers to be encoded as NANs, with additional bits available for tags. [^2] NAN-boxing has entered the mainstream. However, as larger virtual address spaces are supported, such as by Intel’s 5-level page tables, this becomes more complicated.
We anticipate that being standard compliant versus having a finite number of low-level specialisations will remain a nuanced judgment call. The issue of claimed versus actual compliance of language implementations will remain ongoing as standards are periodically revised. For example, the format library was initially unavailable in the most popular C++ implementations compiling to the C++20 standard. Said another way, claimed standard compliance remains untrustworthy.
@@
  • libraries of formalised mathematical proofs (e.g. Lean’s mathlib, Mizar, Isabelle/HOL),
@@
As yet, these areas are not problematic, as there are not multiple variants in heavy use, but it is not unreasonable to anticipate that they may give rise to issues in time.
@@ 8 Portability of Mathematics
The portability of mathematics itself remains an on-going open issue in the portability of computer algebra. Differing definitions occur across mathematics’ many fields. We give a few examples where different authors take opposite stances:
  • Logic and foundations. Do the natural numbers include zero? Does the logic have the law of the excluded middle?
  • Algebra. Rings may be or may not be required to have a multiplicative identity. Lie brackets may be defined as [x,y]=u(xyyx)[x,y]=u(xy-yx) with $u$ being $1$ or $i$. Modules are sometimes implicitly taken to be left modules, but in noncommutative contexts the distinction between left and right multiplication matters.
  • Algebraic geometry. Do divisors on a scheme follow the Weil or Cartier definition? These notions coincide on smooth varieties, but do not on singular varieties.
  • Algebraic topology. Poincaré duality is standard in oriented manifolds, but treatments differ in non-orientable cases.
  • Homological algebra. Boundary operators in simplicial homology can differ by sign conventions, also affecting chain complexes.
  • Graph theory. Do graphs allow loops, that is edges from a vertex to itself? Are complete graphs and edgeless graphs considered to be strongly regular? This affects their classification and algebraic combinatorics.
  • Complex functions. Branch cuts can differ for multivalued functions. For example the principal value of complex logarithm may be defined as $\log z=\log|z|+i\arg z$, where $\arg z\in(-\pi,\pi]$ or $0,2\pi)$, depending on convention. Definitions of Fourier transforms vary on whether the forward or inverse transform has the negated exponent and how the forward and inverse transforms are normalised.
  • Differential geometry. The Riemann curvature tensor differs by a sign depending on convention. This affects derived quantities such as the Ricci tensor and scalar curvature. The Hodge star operator and volume forms depend on the metric signature and orientation.
  • Formal proofs of the algorithms used. Some will view formal proofs as either not especially important or as clearly too challenging to incorporate. Others take the view that without them the output from mathematical software has to be viewed with suspicion. One case where some implementation of Reduce follows the path of caution is by using a library for double precision elementary functions that comes with formal proofs that the results are accurate and correctly rounded to the last bit. It does this perhaps more because this makes cross-platform regression test outputs match precisely than because it believes most of its users care about those low bits! At a higher level those who wish to see the algorithmic components of an algebra system proved properly will need to make their stances as regards the axiom of choice and the Riemann hypothesis clear, since various transformations rely on those.
@@
A second form of impediment to the portability of mathematics is the encoding of mathematical algorithms. Significant tracts of mathematical knowledge are captured in the syntax of computer algebra and proof assistant systems. Examples were given in Section 2 of two approaches to symbolic integration locked in unfrequented systems.
@@ 9 Final Thoughts and Conclusions
In the platform consolidation period, the set of suitable widely used environments simplified. Computer architectures were for the most part using 8-bit bytes, memory was byte addressable, and the Unicode, Ieee 754 and XML standards could be relied upon for the most part for data representation. The posix standard provided common abstractions across operating systems, though bespoke memory management was still widely used. Most systems or stand-alone programs were implemented in Lisp, C, C++ or Fortran, all of which had mature language standards with mostly compliant implementations. Freely available cross-platform implementations meant implementors could rely on one particular implementation (e.g. g++ or clang/llvm). The Intel instruction architecture was dominant, though that is no longer the case. Today, older legacy systems or modern embedded systems may have diverse architectural considerations that can be handled by earlier methods.
We are now in the mathematical consolidation period where the main remaining portability issues are at the levels of mathematical definitions and algorithm encoding. These come into play when it is desired to share mathematical data or computations across different computer algebra systems and mathematical proof assistants. The issue of consistency of mathematical definitions has existed since well before the digital computer era. Seemingly slight differences in the meaning of specific mathematical symbols can make what is an identity or theorem in one system be false in another system. Examples of issues that arise include sign conventions, normalisation factors, branch cut choices for inverse complex functions, and foundational logics. The issue of algorithm encoding is that much of mathematical knowledge is captured in the form of programs written in the application-level programming language of a handful of systems such as Magma, Maple, Mathematica, Reduce and SageMath (for computer algebra) or Mizar, Coq, Isabelle/HOL or Lean (for proof assistants). Here, not only may the details of definitions differ, but the encoding of algorithms is by-and-large completely incompatible.
@@
The second conclusion is that portability is not automatically easy, and the main challenges have altered over time-scales that are relevant to some software. As outlined above, we began in the 1960s with a small number of wildly different machines, but software that was simple by today’s standards. Over the next decades, into the 1990s, many new platforms were introduced and standardisation of languages and interfaces was a work in progress. We have reached a point where there are fewer machines that are relevant again, and much standardisation to help, but software is much more ambitious. There are, however, a variety of portability issues where the originator of the code has not been sloppy, but the world has changed beneath their feet. For example, systems are now concerned with security or protecting intellectual property, resulting in uncontrolled data relocation, unwritable code segments, code signing or license managers. Compilers with aggressive code optimization may exploit all behaviour allowable by the language standard, invalidating previously reasonable assumptions about 2-s complement arithmetic or instruction ordering. So things are sometimes tough again.
@@
The fourth conclusion is that freely available algorithmic code that runs atop a commercial platform may feel locked down if it is complex. Arranging that it is portable to different worlds (either commercial or open source) is a long way from automatic. So the fact that somebody has a paper describing how they can do something you like may or may not help. And if they then cease to be available to support what they had done the progress they made may be lost. So portability there is critical!
The final conclusion is that the people who care about portability are mostly invisible, but rest of the world depends on the work they do. Take, for example, the many many who keep Linux afloat. For (much) smaller projects there is a (much) smaller pool of expertise, but maintenance and porting are very close neighbours. We can not offer magic bullet solutions to either, but feel that highlighting the issues is important.